OpenStudy (anonymous):

Subtract: 5/(x-7) - 4/(3x-11)

OpenStudy (anonymous):

I just need to check my answer :)

OpenStudy (anonymous):

2(2x+2)+3x=11 4x+4+3x=11 7x+4=11 7x+4-4=11-4 7x=7 x=1

OpenStudy (anonymous):

What did you do??

OpenStudy (anonymous):

\[(11x-83)/(3x^2-32x-77)\]

OpenStudy (anonymous):

Thank you :)

OpenStudy (zarkon):

hmmm \[\frac{11x-27}{3x^2-32x+77}\]

OpenStudy (anonymous):

Gahhhh! That's the answer i got, But someone posted 11x-83/3x^2-32x-77......

OpenStudy (zarkon):

\[\frac{5}{x-7} - \frac{4}{3x-11}\] \[=\frac{5(3x-11)-4(x-7)}{(x-7)(3x-11)} \] \[=\frac{15x-55-4x+28}{(x-7)(3x-11)} \] \[=\frac{11x-27}{(x-7)(3x-11)} \] \[=\frac{11x-27}{3x^2-32x+77} \]

OpenStudy (anonymous):

Thank you for clarifying the answer :)

OpenStudy (zarkon):

no problem